If a = −3, b = 4 and c = −5, then by how much is the product abc greater than the sum a + b + c?


1
Expert's answer
2021-05-07T12:25:08-0400

"a = −3, b = 4 \hspace{1mm} and \hspace{1mm} c = −5"

"-3 \times 4 \times 5 = 60"

"(-3)+4-5 = -4"

"60-(-4)"

Answer : 48 Greater than the sum.